[5] Weighted Average Cost of Capital (WACC) (25 points) Apple Inc. has 5,366,166,000 shares traded at a market value of $126 per share, its debt outstanding has an estimated market value of $90,883,140,000. The stock has a beta of 1.193. The expected retum on stocks is 8.78%. The firm is rated AAA and paid 30% of its income as taxes. The risk-free rate is 2.47%, and the yield-to-maturity of Apples bonds is 3 2796. A. What is Apples debt ratio? B. What is Apples equity ratio? C. What is Apples after-tax cost of debt? D. What is Apples investors required return on equity? E. What Apples cost of capital (after-tax WACC)?
Answer #1

Answer:

Shares Outstanding= 5,366,166,000

Market price=$126 per share

Total Equity=126*5,366,166,000=$676,136,916,000

Total Debt=$90,883,140,000

A)

Debt ratio=Total debt/(Total Debt+Total equity)=90,883,140,000/(90,883,140,000+676,136,916,000)

Debt Ratio=0.1185=11.85%

B)

Equity Ratio=Total Equity/(Total Debt+Total equity)=676,136,916,000/(90,883,140,000+676,136,916,000)

Equity Ratio= 0.8815=88.15%

C)

Given Cost of Debt=3.27%

After tax cost of debt=3.27%*(1-Tax rate)=3.27%*(1-30%)=2.289%

D)

Apple's required rate of Return=risk free rate+ Beta*(Expected return of stocks-risk free rate)

Apple's required rate of Return=2.47%+1.193*(8.78%-2.47%)=9.9978%

E)

After Tax WACC=Debt ratio* After tax cost of debt+Equity ratio*Apple's required rate of Return

After Tax WACC=2.289%*11.85%+9.9978%*88.15%=9.0843%